Suppose that the volume of a right circular cylinder is 225 cubic meters and the area of its base is 25 square meters. What is the height of the cylinder?
Answer by Cromlix(4381)

Volume = Pi * r^2 * h
Base = Pi*r^2
Volume = area of base * height
225m^3 = 25m^2 * h
h = 225m^3/25m^2
h = 9m
Height = 9 metres.
Hope this helps.
:-)